OpenStudy (anonymous):

List the first four terms of a geometric sequence with t1 = 4 and tn = -3tn-1.

OpenStudy (harsimran_hs4):

plug n = 2 tn = -3(tn-1) t2 = -3(t1) t2 = -3(4) = -12 for t3 put n = 3 and so on can you proceed now ?